# Bohr-Sommerfeld type surgeries and Disk Potential

### v1

@@ 1.2. Beginnings of a new surgery
A surgery with Bohr-Sommerfeld profile $\Lambda$ switches between two different exact fillings of the disjoint Legendrian $\Lambda_{2}:=\Lambda\sqcup e^{\frac{i\pi}{k}}\Lambda$ where $k$ is the covering number of $\Lambda\rightarrow L_{Y}$. Here we use the $S^{1}$ bundle structure on $Z$. Given an embedded path $c:[-1,1]\rightarrow\mathbb{C}^{*}$ in the punctured complex plane such that $c(-1)=-1,c(1)=1$, we can construct a filling of $\Lambda_{2}$, which we call $\mathcal{L}_{c}$, see Definition 2.9. We call such Lagrangian fillings $\mathcal{L}_{c}$ Bohr-Sommerfeld Profile Lagrangian handles. There are two primitive homotopy classes of such paths - and a Bohr-Sommerfeld profile surgery switches between these two homotopy classes.
@@
Intuitively, we call a Lagrangian $L$ in a symplectic manifold $M$ to have a Bohr-Sommerfeld-Profile handle if under a symplectomorphism from $[a,b]\times Z$ to a codimension 0 submanifold with boundary in $M$, $L$ looks like $\mathcal{L}_{\gamma}$ for some path $\gamma,$ see Definition 2.15. A Bohr-Sommerfeld-profile surgery of such a Lagrangian $L$ is defined as
@@
In Lemma 2.21 we show that if $L$ is a monotone Lagrangian and $\widetilde{\gamma}$ satisfies some properties, then a Bohr-Sommerfeld-profile surgery using $\widetilde{\gamma}$ preserves monotonicity, that is $BSP(L)$ is also monotone. We denote such BSP surgeries as $BSP_{0}(L)$.
@@ 1.3. Acknowledgements
We thank Chris Woodward, Denis Auroux, Mohammed Abouzaid, Georgios Dimitroglou-Rizell, Felix Schlenk and Dylan Cant for valuable discussions. We also thank Amanda Hirschi, Julian Chaidez and Mohan Swaminathan for feedback on an earlier draft. This work was partially supported by the grants NSF DMS 2105417 and NSF DMS 2345030.
@@ 2. Construction of the Bohr-Sommerfeld Profile Surgery
We use certain exact fillings of canonical Bohr-Sommerfeld lifts of monotone Lagrangians as a local model for Lagrangians which allow a generalized surgery. We start with reviewing some preliminaries about constructing Bohr-Sommerfeld lifts of monotone Lagrangians. See [rg19, § 4] for a more in-depth discussion of Bohr-Sommerfeld lifts.
@@ 2.1. Preliminaries
Let $(Y,\omega_{Y})$ be a simply connected, compact, monotone, symplectic manifold with an integral symplectic form. Let $(Z,\lambda)\rightarrow Y$ be a prequantization bundle. Recall that a prequantization over $Y$ is a circle-bundle-with-connection whose curvature is equal to the pullback of the symplectic form $\omega_{Y}$.
@@
In this section whenever we take a parallel transport, it is with respect to the connection induced from $\lambda$.
Let $L_{Y}$ be a monotone Lagrangian submanifold whose minimal Maslov number($N_{L}$) is two.
@@
We assume that the symplectic areas of cycles with boundary on the Lagrangian $L_{Y}$ satisfy,
$[\omega_{Y}]\in\frac{\pi}{k}\operatorname{im}((H_{2}(Y,L_{Y},\mathbb{Z}))\hookrightarrow H_{2}(Y,L_{Y},\mathbb{R})),$
@@
From now on, by convention, we will choose the minimal $k$ which satisfies Assumption 2.2.
@@
Let $E\rightarrow B$ be a circle-bundle with a connection $\lambda$ such that $E$ is flat. Let $P\subset E_{p}$ be a subset of the fiber $E_{p}$ such that it is fixed the holonomy group with base point at $p$, i.e. for any $\gamma:S^{1}\rightarrow B$ with $\gamma(1)=p$, the parallel transport of $P$ along $\gamma$ is $P$. We define the set $\mathfrak{P}$, the parallel transport of $P$ along $B$ to be the set of points $x\in E$ such that $x\in PT_{\gamma}(P)$ for some curve $\gamma:[0,1]\rightarrow B$ with $\gamma(0)=p$.
@@
Since the set $P$ is fixed under holonomy, we have that the set $\mathfrak{P}$ restricted to each fiber is equal to $P$ up to the natural $S^{1}$ action. Now the result follows from the fact that $E$ has no curvature, so locally $\mathfrak{P}$ is a horizontal section over an open neighborhood in $B$. ∎
@@
Let $P_{k}=\left\{e^{i2\frac{l\pi}{k}}\;|\;\;l\in\{1,2,\dots k\}\right\}$ be the set of $k-$th roots of unity. We define the canonical Bohr-Sommerfeld cover $\Lambda\subset Z$ as the manifold obtained by parallel transporting (see Definition 2.3) $P_{k}$ along the Lagrangian $L_{Y}$ after identifying a fiber over a point with $S^{1}$.
Note that $\Lambda$ is a Legendrian submanifold of $Z$. Assumption 2.2 implies that this parallel transport method indeed constructs a closed submanifold of $Z$. See [rg19, § 4.1] for an algebraic construction of $\Lambda$. Note that $\Lambda$ depends on an identification of $S^{1}$ with a fiber over $L_{Y}$, so there is actually a $S^{1}/\mathbb{Z}_{k}\cong S^{1}$ family of $\Lambda$.
The symplectization of $Z$, ×$\mathbb{R}\times Z$ carries a natural ×$\mathbb{R}\times S^{1}\cong\mathbb{C}^{*}$ action induced from the $S^{1}$ action on $Z$ and translation action on the $\mathbb{R}$-component. Thus, we have that ×$\mathbb{R}\times Z\rightarrow Y$ is a $\mathbb{C}^{*}$-bundle. We define a symplectic form on ×$\mathbb{R}\times Z$ as
@@
and any almost complex structure $J_{Y}$ on $Y$ lifts uniquely to a cylindrical almost complex structure $J_{c}$ on ×$\mathbb{R}\times Z$ such that
@@
The almost complex structure $J_{c}$ on ×$\mathbb{R}\times Z$ makes the projection map, $\pi:\mathbb{R}\times Z\rightarrow Y$, a pseudoholomorphic map.
@@
  • there is a $c>0$ such that the $\gamma(t)\in\mathbb{R}$ for $t\in[0,c)\cup(1-c,1].$
@@
Definition 2.8 (Conjugate of admissible curve).
For any admissible curve, a conjugate curve is any curve $\widetilde{\gamma}$ such that
@@
    $[\gamma\#{\widetilde{\gamma}}^{-1}]=1\text{ in }H_{1}(\mathbb{C}^{*})$
@@ 2.2. Bohr-Sommerfeld Profile Lagrangian Handles
Consider the following Legendrian with two connected components, where one component is obtained by rotating a Bohr-Sommerfeld lift by $e^{i\frac{\pi}{k}},$
@@
The truncated symplectization $[r_{1},r_{2}]\times Z$ is an annulus bundle over $Y$ which carries a connection induced by the connection form $\lambda$ on $Z$.
Bohr-Sommerfeld Profile handles are Lagrangian fillings of $\{r_{2}\}\times\Lambda_{2}$ in the truncated symplectization $[r_{1},r_{2}]\times Z$ constructed by parallel transporting. Definition 2.3 of parallel transporting along base can be extended naturally from $S^{1}$ bundles to their associated $\mathbb{C}^{*}$ bundles so we omit the definition.
@@
Given an admissible curve $\gamma:[0,1]\rightarrow A(a,b)\cong[\log a,\log b]$, we define
@@
as the parallel transport of $\gamma^{1/k}$ along the Lagrangian $L_{Y}\hookrightarrow Y$ where we view $[\frac{\log a}{k},\frac{\log b}{k}]\times Z$ as an annulus bundle over $Y$.
Recall that the restriction of the bundle to the Lagrangian $L_{Y}\hookrightarrow Y$ has no curvature, since $d\lambda$ vanishes on $L_{Y}$.
@@
In certain computations, using the identification $A(a,b)\cong[\log a,\log b]\times S^{1}$. We can write an admissible curve using this identification as
@@
Then, one connected component of $\gamma^{1/k}$ is given by $\gamma_{1}(t)=(\frac{1}{k}\gamma_{l}(t),\frac{1}{k}\gamma_{\theta}(t))$.
@@
From the above Remark 2.10 one will notice that the construction of BSP can be naturally generalized when we have the following setup: Let $(Z,\lambda)$ be a contact manifold with a Legendrian $\Lambda$ such that the Reeb flow $\phi_{t}$ is periodic on $\Lambda$ with the same period $T$ i.e. $\phi_{T}|_{\Lambda}=Id$. Further assume that there is a $t<T$ such that $\phi_{t}(\Lambda)=\Lambda$. Then we can use the Reeb flow translates to define two fillings - one by going from time 00 to $t$ and another by going from time 00 to $-(T-t)$.
@@
There are multiple ways to see it. A direct way is to check that the embedding in Remark 2.10 is a Lagrangian embedding. Indeed, we can see that for a fixed $t_{0}$, the embedding $i_{\gamma}(t_{0},\_)$ embeds $\Lambda$ to a phase-shift ( i.e. shifted by the circle action) of the Legendrian $\{\frac{\gamma_{l}(t_{0})}{k}\}\times\Lambda$ in $\{\frac{\gamma_{l}(t_{0})}{k}\}\times Z$. Thus, the result then follows from checking that $(i_{\gamma})_{*}(\partial_{t})$ pairs trivially with $(i_{\gamma})_{*}(v)$ for any $v\in T\Lambda$ which in turn follows
$d(e^{s}\lambda)(R,v)=0$
and
$d(e^{s}\lambda)(\partial_{s},v)=0$
where $v\in T\Lambda$ and $R$ is the Reeb field. ∎
@@
The BSP Lagrangian handle $\mathcal{L}_{\gamma}$ is exact in $([\log a,\log b]\times Z,d(e^{s}\lambda))$. Furthermore, we can choose a function $f_{\gamma}$ on $[0,1]\times\Lambda$ such that $e^{s}\lambda|_{\mathcal{L}_{\gamma}}=d(f_{\gamma}\circ i_{\gamma}^{-1})$ where $i_{\gamma}$ is defined as in Equation 2.2.1. In addition, we can choose $f_{\gamma}$ to depend only on the $[0,1]$ component, i.e. $f_{\gamma}(t,p)=\tilde{f}_{\gamma}(t)$.
@@
Thus, we can define a function $f_{\gamma}$ on $[0,1]\times\Lambda$ by integrating $e^{s}\lambda$ along paths such that
@@
Note that since $e^{i\theta}.\Lambda$ is a Legendrian, we can ensure $f_{\gamma}(t,p)=\tilde{f}_{\gamma}(t)$. Moreover, we have the following explicit formula by choosing $(0,p)\in\{0\}\times\Lambda$ as a base point and integrating paths $\kappa:[0,1]\rightarrow[0,1]\times\Lambda$ such that $\kappa(0)=(0,p),\kappa(1)=(t_{q},q)$. We can homotope $\kappa$ to ensure that $\kappa|_{[0,1/2]}=(2tt_{q},p)$ and $\kappa|_{[1/2,1]}$ is a path in $\{t_{q}\}\times Z$.
@@
Now note that $i_{\gamma}\circ\kappa(t)=\left(\frac{\gamma_{l}(2tt_{q})}{k},e^{i\frac{\gamma_{\theta}(2tt_{q})}{k}}.p\right),$ thus we see that the integral $\int_{\kappa}i_{\gamma}^{*}e^{s}\lambda$ depends solely on $t_{q}$, thus there is a $\tilde{f}$ such that $f_{\gamma}(t,q)=\tilde{f}_{\gamma}(t)$. ∎
@@
Note that the proof above shows that the primitive $\tilde{f}_{\gamma}$ of $i_{\gamma}^{*}e^{s}\Lambda$ depends continuously on the path $\gamma$ and its derivative $\gamma^{\prime}$. By shrinking the interval $[a,b]$ in the embedding $\phi:[a,b]\times Z\rightarrow M$, we can assume that the action of Reeb chords bounding $\Lambda_{2}$ in $\{b\}\times Z$ are rational. From the continuous dependence of $\widetilde{f}_{\gamma}$ on $\gamma$, we see that, by perturbing $\gamma$ appropriately, we can assume $\widetilde{f}_{\gamma}(1)-\widetilde{f}_{\gamma}(0)$ is a rational number. Thus, we can assume any disk with boundary on the closure $\overline{\mathcal{L}_{\gamma}}$ of the BSP handle in the symplectic cutting of $[a,b]\times Z$ at $\{a,b\}\times Z$ has rational symplectic area.

### v2

@@ 1.2. Beginnings of a new surgery
A surgery with Bohr-Sommerfeld profile $\Lambda$ switches between two different exact fillings of the disjoint Legendrian $\Lambda_{2}:=\Lambda\sqcup e^{\frac{i\pi}{k}}\Lambda$ where $k$ is the covering number of $\Lambda\rightarrow L_{Y}$.( Here we are implicitly using the $S^{1}$ bundle structure on $Z$.) Given an embedded path $c:[-1,1]\rightarrow\mathbb{C}^{*}$ in the punctured complex plane such that c(1)=1,c(1)=1c(-1)=-1,c(1)=1, we can construct a filling of $\Lambda_{2}$, which we call $\mathcal{L}_{c}$, see Definition 2.10. We call such Lagrangian fillings $\mathcal{L}_{c}$ Bohr-Sommerfeld Profile Lagrangian handles. There are two primitive homotopy classes of such paths. A Bohr-Sommerfeld profile surgery switches between these two homotopy classes.
@@
Intuitively, we call a Lagrangian $L$ in a symplectic manifold $M$ to have a Bohr-Sommerfeld-Profile handle if under a symplectomorphism from $a,b]\times Z$ to a codimension 0 submanifold with boundary in $M$, $L$ looks like $\mathcal{L}_{\gamma}$ for some path $\gamma,$ see Definition 2.16. A Bohr-Sommerfeld-profile surgery of such a Lagrangian $L$ is defined as
@@
In Lemma 2.24 we show that if $L$ is a monotone Lagrangian and $\widetilde{\gamma}$ is chosen judiciously (see Equation 2.3.3), then a Bohr-Sommerfeld-profile surgery using $\widetilde{\gamma}$ preserves monotonicity, that is $BSP(L)$ is also monotone. We call such surgeries zero-area BSP surgeries and denote them as $BSP_{0}(L)$.
@@ 1.3. Acknowledgements
We thank Chris Woodward, Denis Auroux, Mohammed Abouzaid, Georgios Dimitroglou-Rizell, Felix Schlenk and Dylan Cant for valuable discussions. We thank the anonymous referee whose suggestion helped remove a technical assumption (enough regular disks), making our results significantly stronger. We also thank Amanda Hirschi, Julian Chaidez and Mohan Swaminathan for feedback on an earlier draft. This work was partially supported by the grants NSF DMS 2105417 and NSF DMS 2345030.
@@ 2. Construction of the Bohr-Sommerfeld Profile Surgery
As we have already discussed in the introduction, Bohr-Sommerfeld profile surgeries are obtained from switching between two natural exact fillings of Bohr-Sommerfeld lifts. We start off with reviewing some preliminaries about constructing Bohr-Sommerfeld lifts of monotone Lagrangians. See [rg19, § 4] for a more in-depth discussion of Bohr-Sommerfeld lifts.
@@ 2.1. Preliminaries
Let $(Y,\omega_{Y})$ be a simply connected, compact, monotone, symplectic manifold with an integral symplectic form. Let $(Z,\lambda)\rightarrow Y$ be a prequantization bundle. Recall that a prequantization over $Y$ is a circle-bundle with a connection one-form $\lambda$ whose curvature is equal to the pullback of the symplectic form $\omega_{Y}$.
@@
In this section, whenever we take a parallel transport, it is with respect to the connection induced from $\lambda$.
Let $L_{Y}$ be a monotone Lagrangian submanifold whose minimal Maslov number (which we will denote by $N_{L}$) is two.
@@
We assume that the symplectic areas of cycles with boundary on the Lagrangian $L_{Y}$ satisfy the following,
$\omega_{Y}]\in\frac{\pi}{k}\operatorname{im}((H_{2}(Y,L_{Y},\mathbb{Z}))\hookrightarrow H_{2}(Y,L_{Y},\mathbb{R})),$
@@
Convention 2.3.
From now on, we will choose the minimal positive number $k$ which satisfies Assumption 2.2.
@@
Let $E\rightarrow B$ be a circle-bundle with a connection $\lambda$ such that $E$ is flat. Let $P\subset E_{p}$ be a subset of the fiber $E_{p}$ such that $P$ is fixed by the holonomy group with base point at $p$, i.e. for any $\gamma:S^{1}\rightarrow B$ with $\gamma(1)=p$, the image of $P$ under the parallel transport map along $\gamma$ is $P$ itself. We define the set $\mathfrak{P}$, the parallel transport of $P$ along $B$ to be the set of points $x\in E$ such that $x\in PT_{\gamma}(P)$ for some curve $\gamma:[0,1]\rightarrow B$ with $\gamma(0)=p$.
@@
Since the set $P$ is fixed under holonomy, we have that the set $\mathfrak{P}$ restricted to each fiber is equal to $P$ up to the natural $S^{1}$ action. Now the result directly follows from the fact that $E$ has no curvature, so locally $\mathfrak{P}$ is a horizontal section over an open neighborhood in $B$. ∎
@@
Let $P_{k}=\left\{e^{i\frac{2\pi}{k}l}\;|\;\;l\in\{1,2,\dots k\}\right\}$ be the set of $k$-th roots of unity. We define the canonical Bohr-Sommerfeld cover $\Lambda\subset Z$ as the manifold obtained by parallel transporting (see Definition 2.4) $P_{k}$ along the Lagrangian $L_{Y}$.
Assumption 2.2 and Lemma 2.5 imply that $\Lambda$ indeed is a submanifold of $Z$. Moreover, $\Lambda$ is a Legendrian submanifold of $(Z,\lambda)$. See [rg19, § 4.1] for an algebraic-topology flavoured construction of $\Lambda$. Note that $\Lambda$ depends on an identification of $S^{1}$ with a fiber over $L_{Y}$, so there is actually a $S^{1}/\mathbb{Z}_{k}\cong S^{1}$ family of $\Lambda$.
The symplectization of $Z$, $\mathbb{R}\times Z$ carries a natural $\mathbb{R}\times S^{1}\cong\mathbb{C}^{*}$ action induced from the $S^{1}$ action on $Z$ and translation action on the $\mathbb{R}$-component. Thus, we can view $\mathbb{R}\times Z\rightarrow Y$ as a $\mathbb{C}^{*}$-bundle. We define a symplectic form on $\mathbb{R}\times Z$ as
@@
and any almost complex structure $J_{Y}$ on $Y$ lifts uniquely to a cylindrical almost complex structure $J_{c}$ on $\mathbb{R}\times Z$ such that
@@
The almost complex structure $J_{c}$ on $\mathbb{R}\times Z$ makes the projection map, $\pi:\mathbb{R}\times Z\rightarrow Y$, a pseudoholomorphic map.
@@
  • there is a $c>0$ such that $\gamma(t)\in\mathbb{R}$ for $t\in[0,c)\cup(1-c,1].$
@@
Definition 2.9 (Conjugate of an admissible curve).
For any admissible curve $\gamma$, a conjugate of that curve is any curve $\widetilde{\gamma}$ such that
@@
    $\gamma\#{\widetilde{\gamma}}^{-1}]=1\text{ in }H_{1}(\mathbb{C}^{*})$
@@ 2.2. Bohr-Sommerfeld Profile Lagrangian Handles
We will construct Bohr-Sommerfeld profile surgeries by switching between fillings of the ‘double’ Bohr-Sommerfeld lift $\Lambda_{2}$. This Legendrian has two connected components,
@@
In $\Lambda_{2}$, the connected components are related by a ‘phase-shift’ i.e. action of some element of $S^{1}$, of the other.
@@
We now proceed towards constructing fillings of the Legendrian $\Lambda_{2}$. The truncated symplectization $r_{1},r_{2}]\times Z$ is an annulus bundle over $Y$ which carries a connection induced by the connection form $\lambda$ on $Z$.
Bohr-Sommerfeld Profile handles are Lagrangian fillings of $\{r_{2}\}\times\Lambda_{2}$ in the truncated symplectization $r_{1},r_{2}]\times Z$ constructed by parallel transporting. Definition 2.4 of parallel transporting along base can be extended naturally from $S^{1}$ bundles to their associated $\mathbb{C}^{*}$ bundles so we omit the definition. In certain computations regarding $\mathcal{L}_{\gamma}$, using the identification $A(a,b):=[\log a,\log b]\times S^{1}$ turns out to provide a clearer picture.
@@
Given an admissible curve $\gamma:[0,1]\rightarrow A(a,b)$, we define
@@
to be the parallel transport of $\gamma^{1/k}$ along the Lagrangian $L_{Y}\hookrightarrow Y$ where we view $\frac{\log a}{k},\frac{\log b}{k}]\times Z$ as an annulus bundle over $Y$.
Recall that the restriction of the bundle $Z$ to the Lagrangian $L_{Y}\hookrightarrow Y$ has no curvature, since $d\lambda$ vanishes on $L_{Y}$.
@@
We can write an admissible curve using this identification as
@@
Then, one connected component of $\gamma^{1/k}$ is given by $\gamma_{1}(t)=(\frac{1}{k}\gamma_{l}(t),\frac{1}{k}\gamma_{\theta}(t))$. The other components are given by similar formulae where the $S^{1}$ co-ordinate is shifted by a constant angle.
@@
From Remark 2.11 one will notice that the construction of BSP can be naturally generalized when we have the following setup: Let $(Z,\lambda)$ be a contact manifold with a Legendrian $\Lambda$ such that the Reeb flow $\phi_{t}$ is periodic on $\Lambda$ with the same period $T$ i.e. $\phi_{T}|_{\Lambda}=Id$. Further assume that there is a $t<T$ such that $\phi_{t}(\Lambda)=\Lambda$. Then we can use the Reeb flow to define two fillings - one by going from time 0 to $t$ and another by going from time 0 to (Tt)-(T-t).
@@
There are multiple ways to see it. A direct way is to check that the embedding in Remark 2.11 is a Lagrangian embedding. Indeed, we can see that for a fixed $t_{0}$, the embedding $i_{\gamma}(t_{0},\_)$ embeds $\Lambda$ to a phase-shift ( i.e. shifted by the circle action) of the Legendrian $\{\frac{\gamma_{l}(t_{0})}{k}\}\times\Lambda$ in $\{\frac{\gamma_{l}(t_{0})}{k}\}\times Z$. Notice that we have
$d(e^{s}\lambda)(R,v)=0,$
$d(e^{s}\lambda)(\partial_{s},v)=0,$
where $v\in T\Lambda$ and $R$ is the Reeb field. Thus, the result then follows from the fact that $(i_{\gamma})_{*}(\partial_{t})$ pairs trivially with $(i_{\gamma})_{*}(v)$ for any $v\in T\Lambda$. ∎
@@
The BSP Lagrangian handle $\mathcal{L}_{\gamma}$ is exact in $([\log a,\log b]\times Z,d(e^{s}\lambda))$. Furthermore, we can choose a function $f_{\gamma}$ on $0,1]\times\Lambda$ such that $e^{s}\lambda|_{\mathcal{L}_{\gamma}}=d(f_{\gamma}\circ i_{\gamma}^{-1})$ where $i_{\gamma}$ is defined as in Equation 2.2.1. In addition, we can choose $f_{\gamma}$ to depend only on the $[0,1]$ component, i.e. $f_{\gamma}(t,p)=\tilde{f}_{\gamma}(t)$.
@@
Thus, we can define a function $f_{\gamma}$ on $0,1]\times\Lambda$ by integrating $e^{s}\lambda$ along paths such that
@@
Note that since $e^{i\theta}.\Lambda$ is a Legendrian, we can ensure $f_{\gamma}(t,p)=\tilde{f}_{\gamma}(t)$. Moreover, we have the following explicit formula by choosing $(0,p)\in\{0\}\times\Lambda$ as a base point and integrating paths $\kappa:[0,1]\rightarrow[0,1]\times\Lambda$ such that $\kappa(0)=(0,p),\kappa(1)=(s_{q},q)$. We can homotope $\kappa$ to ensure that $\kappa|_{[0,1/2]}=(2ts_{q},p)$ and $\kappa|_{[1/2,1]}$ is a path in $\{s_{q}\}\times Z$.
@@
Now note that $i_{\gamma}\circ\kappa(t)=\left(\frac{\gamma_{l}(2ts_{q})}{k},e^{i\frac{\gamma_{\theta}(2ts_{q})}{k}}.p\right),$ thus we see that the integral $\int_{\kappa}i_{\gamma}^{*}e^{s}\lambda$ depends solely on $s_{q}$, thus there is a $\tilde{f}$ such that $f_{\gamma}(t,q)=\tilde{f}_{\gamma}(t)$. ∎
@@
Note that the proof above shows that the primitive $\tilde{f}_{\gamma}$ of $i_{\gamma}^{*}e^{s}\Lambda$ depends continuously on the path $\gamma$ and its derivative $\gamma^{\prime}$. By shrinking the interval $[a,b]$ in the embedding $\phi:[a,b]\times Z\rightarrow M$, we can assume that the action of Reeb chords bounding $\Lambda_{2}$ in $\{b\}\times Z$ are rational. From the continuous dependence of $\widetilde{f}_{\gamma}$ on $\gamma$, we see that, by perturbing $\gamma$ appropriately, we can assume $\widetilde{f}_{\gamma}(1)-\widetilde{f}_{\gamma}(0)$ is a rational number. Thus, we can assume any disk with boundary on the closure $\overline{\mathcal{L}_{\gamma}}$ of the BSP handle in the symplectic cutting of $a,b]\times Z$ at $\{a,b\}\times Z$ has rational symplectic area. These rationality results will be pertinent while constructing broken divisors to make regularize moduli space of SFT buildings using domain dependent perturbations.